# On the Potential of Microtubules for Scalable Quantum Computation

### v1

@@ II Solitonic effects in Microtubules and observable biological functions
[…]
@@ III Microtubular networks as logic gates
Figure 7: A MT arrangement in cell as a ‘logic’ XOR gate. Left panel: the biological arrangement, a group of MTs and their MAPs, reproduced from  [1]. Right panel: a conventional XOR gate for comparison.
[…]
@@ IV Scalable Quantum Computation in Microtubules
To this end, we first remark that above we have treated the solitons on a MT as classical solutions, and the XOR gate role of MT arrangements was construed according to classical physics and computation. As in our previously published work  [3, 4], these solitons are treated as macroscopic quantum coherent states, or at most as (not completely decohered, minimum entropy) pointer states [16], which survive long enough so that processes such as energy transmission (and by extension memory storage) along MTs of length of a few $\mu$m are allowed to take place. In our approach, this can happen under fine-tuned conditions [3], that defend the system from environmental decoherence losses. The dipole-dipole interactions between the tubulin dimers and ordered-water dipole quanta [51] in the neighborhood of the (hydrophobic) walls of the MT tubulin dimers are crucial in accomplishing this.
@@ IV.2 Microtubules as Quantum Computers: Detailed Scheme
$|\mathcal{A}\rangle=|\alpha\beta\alpha\beta\rangle-|\beta\alpha\beta\alpha\rangle\,$
@@
These quantum states are quantum entangled (“wired dissipationlessly”) during the time interval (14) (measured from the moment of the action of an external stimulus to the MT), in much the same way as the bilin molecules of the algae antennae. In algae, such a coherent wiring / yoking / quantum entanglement occurs over distances of 25 nm, and the decoherence time is a few hundreds fs. This is evidently sufficient time for the algae molecule to quantum-compute the optimal path for signal transmission over distances covering half of the extent of the algae light harvesting antenna. In the case of the MT, the distances, over which entanglement is expected to survive, are not restricted only to the fundamental cell depicted in fig. 5, but, over the entire MT of lengths of $\mathcal{O}(\mu\rm m)$. This feature is a consequence of the strong dipole-dipole interactions between nearest tubulin neighbours, and also the ordered-water dipole quanta [51, 3] and the tubulin dimers themselves, which provide a stronger isolation than in the case of algae, thus leading to much longer coherence times (14). Indeed, the reader should notice that, if the snoidal waves propagate with a velocity of order at most 155 m/s, as discussed after (22), then signals are transduced across a micron ($\mu$m)-long MT in times $\mathcal{O}(10^{-8})$ s, which lies comfortably in the aforementioned decoherence time interval of the QED-cavity model of MT [3].
@@
Another question is whether, via dissipation-free information and energy transmission, quantum computation could be sustained. Here is where the analogy with the ”quantum wiring” encountered in marine algae, as discussed in the introduction [20, 21] serves as an appropriate perspective. There is a direct analogy of the MT system with the algae system, provided we account for the effects of ordered water, which, while traditionally underappreciated has strong enough dipole-dipole interactions [3], to provide the necessary environmental isolation so that quantum effects in tubulin dimers experience decoherence times (14), of order $\mu$s -clearly potentially relevant to many other processes taking place in a living cell as well.
@@ V.1 Testing the Cavity-MT Model: Rabi-Splitting
$\Omega=\omega_{0}-\frac{\Delta}{2}\pm\frac{|\Delta|}{2}\,\Big{(}1+\frac{4\,\lambda^{2}\,\rm N}{\Delta^{2}}\Big{)}^{1/2}\,,$ (31)
where $\lambda$ is the so-called Rabi coupling, pertaining to the atom-photon interactions, viewing the atoms as spin-1/2 two-state quantum systems, and $\rm N$ is the number of atoms inside the cavity. Representing each tubulin in a MT as a two-state ($\alpha$, $\beta$ conformations) ”atom” in a cavity near the dimer walls, induced by the strong dipole-dipole interactions between the dimers and the ordered-water molecules, we covered in detail in [3] showing that a Rabi-like phenomenon similar to that in (31) could be used to characterize these entities at ambient temperatures, during the short decoherence time (14), which is itself linked to the magnitude of the Rabi coupling between the tubulin dimers and the coherent modes of the water-dipole quanta [51] which play the role of ”photons” in this formulation  [3].
@@
where $d_{\rm dimer}$ represents the matrix element of the electric dipole of a single dimer, associated with the transition from the $\alpha$ to the $\beta$ conformations (for relevant physical parameter values see Table I), which provide the binary nature of the dimer quantum state in our framework, as discussed above. The quantity $E_{\rm ow}$ represents a typical root-mean-square (r.m.s.) value of the amplitude of a coherent dipole field mode in the ordered water [51]. As a crude estimate, taking into account that, in contrast to the dimers which lie in the surface of a MT, the ordered-water dipole quanta exist in the entire water-interior of the MT, we may borrow relevant formulae from quantum optics of dielectric cavities, to represent
@@
As discussed in [3] this frequency is in the range of the upper bound of the assumed range of frequencies of quantum oscillations of the tubulin dimers viewed as two-state quantum systems ($\alpha$, $\beta$ conformations), before collapse, as per the original analysis of [15], and in [64]. Our study of MT as QED cavity models in [3] and here assumes this upper bound, and thus, in this case, the dominant cavity mode and the dimer system are almost in resonance, with a detuning satisfying $\Delta\equiv\omega_{0}-\omega_{c}\ll\omega_{0}$, where $\omega_{0}\sim 10^{12}~{}\rm Hz$ the dimer conformational quantum oscillations. In [3] when estimating the relevant number of dimers (“atoms”) in the cavity, we restricted ourselves to one-dimensional solitons formed along the protofilament of an MT. In the current work, we consider helicoidal solitons on the MT surface [7], and therefore we need to consider the number of dimers in the entire MT. For a typical moderately long MT, of length or order of a micron ($\mu$m), consisting of 13 protofilaments, there are
$\mathcal{N}={\mathcal{O}}(10^{3})\,,$ (35)
@@
where we took into account that the vacuum permittivity is $\epsilon_{0}=55.26\,e^{2}\,{\rm eV}^{-1}\,(\mu{\rm m})^{-1}$. On the other hand, for an MT of length 25 $\mu\rm m$, we obtain $E_{\rm ow}\sim 3.5\times 10^{4}\,\frac{\rm V}{\rm m}$.
@@
$\lambda_{\rm MT}=\sqrt{\mathcal{N}}\,\lambda_{0}\sim 1.2\times 10^{12}~{}\rm Hz\,.$ (37)
A typical detuning considered in [3] is such that $|\Delta|/\lambda_{0}\sim{\mathcal{O}}(10^{2})$, implying a situation that $\lambda_{0}^{2}\,\mathcal{N}\ll|\Delta|$, which implies an approximate Rabi splitting for the case of a typical MT arrangement in the biological cells [3]:
$\Omega\simeq\omega_{0}\pm\frac{{\mathcal{N}}\,\lambda_{0}}{|\Delta|^{2}}+{\mathcal{O}}(\Delta)\,.$ (38)
We remark at this point that, in typical Rabi splitting situations in quantum optics it is assumed that practically no energy exchange takes place between atoms and cavity modes. In the MT case, this may be guaranteed from the fact that, since the dominant interactions between ordered-water coherent dipole models (“cavity” modes in this analogue) and dimer-dipole quanta (“atoms”) attenuate with the cubic power of the distance between then, the only dominant interactions are near the walls of the MT, thereby implying that the bulk of the cavity modes (viewing the entire MT as an isolated cavity) does not exchange significant amount of energy with the dimers. This lead us in [3] to assume that the main reason of decoherence is the leakage of dipole quanta from the MT interior to the environment, which lead to the principal estimate of decoherence time (14), which is much larger than the decoherence time (29) of [53], based on individual dimer dipole-environment interactions, outside the cavity model.
@@ V.2 Probing quantum coherence and environmental entanglement of individual tubulin dimers
Table 2: Compiled physical, chemical, and quantum parameters relevant to the study of microtubules as potential substrates for quantum information processing. All values correspond to physiological (’ambient”) conditions we have defined as ($37^{o}\,{\rm C}$, pH 7.2, salinity $\sim 150$ mM), and where applicable, values are normalized or expressed in units of thermal energy ($kT$), GTP hydrolysis energy, or standard quantum energies. experimentally) Dielectric constant of tubulin and MTs (high-frequency): $\kappa=8.41$. High-frequency polarizability of tubulin: $\alpha=2.1\times 10^{-33}\,\rm C\,m^{2}/V.$ Sources include both theoretical predictions and experimental results across cryogenic and biological regimes. We also note the finding of electrical oscillations of bundles of microtubules in the brain [37].
@@
Inner lumen diameter 1414141415151515 nm
@@
Typical microtubule length (neuronal) 1111100100100100 µm Varies with cell type[33]
Cryptochrome decoherence time $>10^{-5}$ s Predicted, bird magnetoreception [77]
Photosystem I coherence time 300300300300800800800800 fs [78]
Light harvesting complex (FMO) coherence $>300$ fs Room temp evidence [78]
@@
Soliton propagation speed (microtubules) 222220202020 m/s Nonlinear excitations [33]
Time for soliton to travel 1 µm 50505050500500500500 ns $\tau=\frac{1\,\mu m}{v}$
@@
These values are 22223333 orders of magnitude smaller than $k_{B}T$, implying that such weak fields alone are insufficient to induce dipole state transitions in thermal equilibrium unless enhanced local fields are used such as those created by defects or nanostructures at a conducting surface. Even without such assistance (which can increase the field strength by several orders of magnitude locally), in the structured environment of the microtubule interior—particularly when interacting with ordered water and neighbouring dipoles—these couplings may contribute to the initiation or modulation of coherent quantum dynamics. In this sense, field-driven transitions may act as subtle biasing agents rather than direct triggers of quantum state evolution.
@@
[…]
@@
By varying the plasmonic path length (200–800 nm), it should be possible to measure decoherence times and assess the persistence of entanglement post-interaction, thereby testing the viability of protein-based “bioqubits” in biological quantum computing contexts.

### v2

@@ II Solitonic effects in Microtubules and observable biological functions
Apart from kinks, another interesting solution for the case $\kappa=0$ are Spikes. The latter are excitations of the ground state. Estimates of the energy carried by a spike have been provided in the analysis of [7]. The electric field produced by the spike can be estimated as [7], $\Delta E_{z}=E_{z}^{\rm max}(u_{sp}^{2}-u_{g}^{2})^{2}$, where $E_{z}^{\rm max}=\frac{Jg_{1}}{S}$ is the maximum value of the electric field due to the permanent dipoles, which is reached when all dipoles are aligned along the MT (in which $u_{g}=1$). The maximum value of the electric field produced by spike $\Delta E_{z}\leq\Delta E_{z}^{\max}$, has been estimated in [7] as $\Delta E_{z}^{\max}=E_{z}^{\rm max}(1-u_{g}^{2})^{2}={E_{z}^{\rm max}}\cos^{4}\Theta_{0}\leq E_{z}^{\rm max}$, where $\Theta_{0}$ denotes the angle between the permanent dipole and axis orthogonal to the surface of the MT. Notice that the maximum magnitude of the electric field produced by spike is bounded by $E_{z}^{\rm max}$. As discussed in the literature [35], and mentioned above, in the ground state the orientation of the dipoles with respect to the surface of the MT are $\Theta_{0}\approx 29^{\,\rm o}$. Taking into account data from [7], one can arrive at the following estimation for the electric field produced by the spike: $\Delta E_{z}^{\max}\approx 0.6{E_{z}^{\rm max}}$. To evaluate $E_{z}^{\rm max}$, we use data available for the electric field inside of the MT: $E_{z}\sim 10^{5}\div 10^{8}\,\,\rm V/m$ [32]. Then, we arrive at the following estimate for the electric field produced by the spike: $\Delta E_{z}^{\rm max}\lesssim\,0.6\cdot(10^{5}\div 10^{8})\,\,\rm V/m~.$
@@ III Microtubular networks as logic gates
Figure 8: A MT arrangement in cell as a ‘logic’ XOR gate. Left panel: the biological arrangement, a group of MTs and their MAPs, reproduced from  [1]. The MT (a) acts as an “input”, whilst (b) acts as an “output”, with (c) denoting a MAP transmitting a soliton, while (d) represents a “quite” map. Right panel: a conventional XOR gate for comparison. A and B act as input, while Q is the output. For a XOR, the truth table reads: A=0, B=0, then Q= 0; If any of A,B is 1 and the other 0, then Q=1, while if both A and B are 1, then Q=0.
In addition to efficient energy transport, the presence of solitonic structures in MT may imply their role as biological logic gates, as proposed initially in [4], and elaborated further, from a quantum computational viewpoint in [1] (See fig. 8, left panel). Although MT do not themselves branch, the analogue of a ‘logic’ XOR gate (see Fig. 8, right panel) by MT arrangements in cells has been proposed with microtubule associated proteins (MAPs) that connect the various MTs in a network given the role of information storage  [17, 33, 18]. Once a soliton is formed along one MT, an “active MAP” (yellow colour in left panel of Fig. 8) can transport it from one MT to another. In the MT arrangement depicted on the left panel of Fig. 8, an XOR logic gate can be realized provided the “0” entry is represented by the absence of a soliton and the “1” entry by the presence of a soliton. In this arrangement, MT (a) acts as the “Input” MT, whilst MT (b) is the “Output” MT. (c) is a MAP transmitting a soliton, while (d) represents a “quiet” MAP (green coloured MAPs). MT (a) has two solitons travelling (yellow colour), encountering two MAPs (yellow coloured MAPs) that transmit both solitons to MT (b). The solitons would arrive out of phase at MT (b) and cancel each other out. The truth table for XOR reads: 0,$0$; $0\rightarrow 0;\,0,1\rightarrow 1;\,1,0\rightarrow 1;\,1,1\rightarrow$ 0\,, and in this case is realized by MTs if the MAPs are arranged in such a way that each can transmit a soliton independently but if they both transmit, then the two solitons cancel each other out. We stress that the existence of snoidal waves is crucial for the behaviour of MT as logical gates, since for such solitonic structures, the out-of-phase cancellation 1,1 $\rightarrow\,$ 0 exists automatically.
@@ IV Scalable Quantum Computation in Microtubules
[…]
@@ IV.2 Microtubules as Quantum Computers: Detailed Scheme
$\mathcal{A}\rangle=|\alpha\beta\alpha\beta\rangle-|\beta\alpha\beta\alpha\rangle\,.$ (31)
@@
[…]
@@
[…]
@@ V.1 Testing the Cavity-MT Model: Rabi-Splitting
$\Omega_{\pm}=\omega_{0}-\frac{\Delta}{2}\pm\frac{1}{2}\,\Big(\Delta^{2}+4\,\lambda^{2}\,\mathcal{N}\Big)^{1/2}\,$, (32)
where $\lambda$ is the so-called Rabi coupling, pertaining to the atom-photon interactions, viewing the atoms as spin-1/2 two-state quantum systems, and $\rm N$ is the number of atoms inside the cavity.
Representing each tubulin in a MT as a two-state ($\alpha$, $\beta$ conformations) “atom” in a cavity near the dimer walls, induced by the strong dipole-dipole interactions between the dimers and the ordered-water molecules, we covered in detail in [3] showing that a Rabi-like phenomenon similar to that in (32) could be used to characterize these entities at ambient temperatures, during the short decoherence time (14), which is itself linked to the magnitude of the Rabi coupling between the tubulin dimers and the coherent modes of the water-dipole quanta [51] which play the role of ”photons” in this formulation  [3].
@@
where $d_{\rm dimer}$ represents the matrix element of the electric dipole of a single dimer, associated with the transition from the $\alpha$ to the $\beta$ conformations (for relevant physical parameter values see Table 3), which provide the binary nature of the dimer quantum state in our framework, as discussed above. The quantity $E_{\rm ow}$ represents a typical root-mean-square (r.m.s.) value of the amplitude of a coherent dipole field mode in the ordered water [51]. As a crude estimate, taking into account that, in contrast to the dimers which lie in the surface of a MT, the ordered-water dipole quanta exist in the entire water-interior of the MT, we may borrow relevant formulae from quantum optics of dielectric cavities, to represent
@@
As discussed in [3] this frequency is in the range of the upper bound of the assumed range of frequencies of quantum oscillations of the tubulin dimers viewed as two-state quantum systems ($\alpha$, $\beta$ conformations), before collapse, as per the original analysis of [15], and in [64]. Our study of MT as QED cavity models in [3] assumed this upper bound, and thus, in this case, the dominant cavity mode and the dimer system were almost in resonance, with a detuning satisfying $\Delta\equiv\omega_{0}-\omega_{c}\ll\omega_{0}$, where $\omega_{0}$ is the frequency of the dimer conformational quantum oscillations. In [3] when estimating the relevant number of dimers (“atoms”) in the cavity, we restricted ourselves to one-dimensional solitons formed along the protofilament of an MT.
[…]
$\mathcal{N}={\mathcal{O}}(10^{2})\,$, (36)
@@
where we took into account that the vacuum permittivity is $\epsilon_{0}=55.26\,e^{2}\,{\rm eV}^{-1}\,(\mu{\rm m})^{-1}$. On the other hand, for an MT of length 25 $\mu\rm m$, which we shall make use of in the subsequent analysis for concreteness, we obtain
$\mathcal{N}\sim 3\times 10^{3}\,$, (38)
and
$E_{\rm ow}\sim 3.6\times 10^{4}\,\frac{\rm V}{\rm m}.$ (39)
Taking into account that each dimer has a mobile charge (positive or negative) 36ee, with $e$ the positron charge, with electric dipole moment  [1]: $d_{\rm dimer}\sim 3\times 10^{-18}$ Cb $\times$ Angström. This implies a Rabi splitting for the individual dimers (33) of order
$\lambda_{0}\sim 1.0\times 10^{11}~\rm Hz\,$, (40)
and a total Rabi splitting for the entire MT, consisting of 13 protofilaments, of order [63]
$\lambda_{\rm MT}=\sqrt{\mathcal{N}}\,\lambda_{0}\sim 5.5\times 10^{12}~\rm Hz\,.$ (41)
Moreover, the detuning $\Delta$ is a phenomenological parameter in the current approach, which goes beyond the chain approximation for the MT model presented in [3], where only the case of a resonant cavity was considered. The two branches of the function $\Omega(\Delta)$ (32) are plotted in fig. 9 (and, for a bigger range of values of $\Delta$ in fig. 10), which expresses the Rabi slitting as a function of a general detuning parameter $\Delta$. Such plots will hopefully prove useful in guiding the relevant experimental searches. The allowed range of $\Delta$ is determined by the requirement of positive absorption frequencies $\Omega_{\pm}>0$, which is depicted in the figures.
We appreciate here that an absorption spectrum of MTs in solution or on a surface would be expected. In practice, intact microtubules (MTs) in suspension do not yield a unique, time-invariant UV–visible “absorption spectrum.” Proteins—including tubulin and the same protein when polymerized into MTs exhibit a broad near-UV maximum near 280 nm arising from aromatic residues (Trp/Tyr) [65, 66, 67]. This intrinsic peak position is not expected to shift with filament length per se; yet upon polymerization the sample’s strong elastic light scattering (“turbidity”), generally swamps weak absorbance features and is proportional mainly to the mass of polymerized material rather than the exact length distribution [69, 68]. Moreover, the hallmark dynamic instability of MTs continuously alters the polymer mass and length distribution on experimental timescales, further causing baseline drift and inhomogeneous broadening that preclude a stable spectrum for freely suspended MTs [70]. For these reasons, the field’s standard quantitative readouts in bulk are turbidimetric (350 nm) rather than true absorbance spectra [69], and optical/electrical characterization that avoids turbidity typically relies on well-defined tubulin solutions (e.g., refractometry/SPR for dielectric properties) or immobilized/stabilized MT formats [71]. We therefore have cited representative spectra and parameters for tubulin in the literature that apply in solution (which report the underlying chromophores) and point to stabilized/immobilized MT measurements or dielectric/impedance spectroscopy for polymer-state properties that are not confounded by scattering [72].
@@
[…]
Some of our results and parameters assumed are tabulated in Tables 2 and 3.
Quantity Estimate / Formula
[…]
Resonant split $\Omega_{\rm split}\approx 2\lambda_{0}\sqrt{N}$ (limiting form of Eq. (32)), which lies in the GHz–THz range for SPR-like fields $10^{4}$$10^{5}$ V/m.
Decoherence window $t_{\rm decoh}\sim 10^{-6}$ s (from Eq. (14)), with dielectric sensitivity shown in Fig. 11.
Table 2: Quantitative benchmarks for MTs referenced in Section V.
@@
We remark at this point that, in typical Rabi splitting situations in quantum optics it is assumed that practically no energy exchange takes place between atoms and cavity modes. In the MT case, this may be guaranteed from the fact that, since the dominant interactions between ordered-water coherent dipole models (“cavity” modes in this analogue) and dimer-dipole quanta (“atoms”) attenuate with the cubic power of the distance between then, the only dominant interactions are near the walls of the MT, thereby implying that the bulk of the cavity modes (viewing the entire MT as an isolated cavity) does not exchange significant amount of energy with the dimers. This lead us in [3] to assume that the main reason of decoherence is the leakage of dipole quanta from the MT interior to the environment, which lead to the principal estimate of decoherence time (14). In terms of the microscopic parameters of the model defined above, the decoherence time reads [3]:
$t_{\rm ow-decoh}=\frac{T_{r}}{2n{\cal N}{\rm sin}^{2}\left(\frac{{\cal N}n\lambda_{0}^{2}t}{\Delta}\right)}\,$, (42)
where $n$ is the average number of oscillation quanta in a coherent mode of dipole, taken in [3] to lie in the range $n=\mathcal{O}(1-10)$. The time $t$ appearing in (42) represents the ‘time’ of interaction of the dimer system with the dipole quanta, which in [3] has been taken to be the average life-time of a coherent dipole-quantum state. In the super-radiance model for the ordered water of [64], which is used in [3] to arrive at (42), this is estimated as
$t\sim\frac{c\hbar^{2}V}{4\pi d_{ej}^{2}\,\Delta E_{\rm ow}^{\rm principal}\,N_{w}\,L}$ (43)
with $d_{ej}$ the electric dipole moment of a water molecule, $L$ the length of the MT, and $N_{w}$ the number of water molecules in the volume $V$ of the MT. We remind the reader at this point that the quantity $\Delta E_{\rm ow}^{\rm principal}$ denotes the energy difference between the two principal energy eigenstates of the water molecule, which are assumed in [64] and [3] to play the dominant rôle in the interaction with the (quantized) electromagnetic radiation field. For typical values of the parameters for moderately long MT, $L\sim 10^{-6}~{\rm m}$, $N_{w}\sim 10^{8}$, a typical value of $t$ is: $t\sim 10^{-4}~{\rm sec}$.
In (42), we took into account that in the QED-cavity model for MT [3], the dominant (dimer)-(dipole quanta) coupling occurs for ordered-water ‘cavity’ modes which are almost at resonance with the dimer oscillators slightly detuned by $\Delta:~\lambda_{0}/\Delta<<1$. Moreover, in [3] we assumed that the time scale $T_{r}$ over which a cavity MT dissipates its energy is of similar order as $t$:
$T_{r}\sim t\,.$ (44)
Under these approximations, the final expression for the decoherence time (42) is:
$t_{\rm ow-decoh}\sim\frac{2\pi\,\Delta^{2}\,d_{ej}^{2}\,\Delta E_{\rm ow}^{\rm principal}\,N_{w}\,L}{n^{3}\,\mathcal{N}^{3}\,\lambda_{0}^{4}\,c\,\hbar^{2}\,V}=\frac{\Delta^{2}\,d_{ej}^{2}\,\Delta E_{\rm ow}^{\rm principal}\,N_{w}\,L\,V}{2\pi\,n^{3}\,\mathcal{N}^{3}\,d_{\rm dimer}^{4}\,\omega_{c}^{2}\,c\,}\,(\epsilon_{0}\,\varepsilon)^{2}\equiv\mathcal{A}\,\varepsilon^{2}\,$, (45)
where, in arriving at the last equality on the right-hand side, we made use of (33) and (34).
Figure 11: The decoherence time (45) in the QED-Cavity model of MT of [3], as a function of the (dimensionless) dielectric constant $\varepsilon$ of the ordered-water medium. The increase of the decoherence time with increasing dielectric constant $\varepsilon$ is easily understood by the fact that an increase in $\varepsilon$ implies a reduction in the strength of the Rabi interaction coupling (33) between dimers and water dipole quanta, which in turn implies a weaker coupling of the system with its environment, and thus a reduction in the losses through the dimer walls. For the parameters of the model in [3], for $\varepsilon=80$ the resulting decoherence time is estimated to lie in the range (14), i.e. $t_{\rm ow-decoh}={\mathcal{O}}(10^{-7}-10^{-6})$ s. We also note, for completeness, that there is a slight dependence of the dielectric constant of water $\varepsilon_{w}$ with temperature T: $\varepsilon_{w}\simeq 80$ at T=20 oC (Celcius), $\varepsilon_{w}\simeq 73.151$ at T=40 oC and $\varepsilon_{w}\simeq 55.72$ at T=100 oC. The reader should always have in mind, though, that in in vivo biological systems, which may not always be in thermal equilibrium, the concept of temperature might be subtle, so plots like this in the figure are mainly used to guide in vitro situations.
[…]
@@ V.2 Probing quantum coherence and environmental entanglement of individual tubulin dimers
Table 4: Compiled physical, chemical, and quantum parameters relevant to the study of microtubules as potential substrates for quantum information processing. All values correspond to physiological (’ambient”) conditions we have defined as ($37^{o}\,{\rm C}$, pH 7.2, salinity $\sim 150$ mM), and where applicable, values are normalized or expressed in units of thermal energy (kTkT), GTP hydrolysis energy, or standard quantum energies. experimentally) Dielectric constant of tubulin and MTs (high-frequency): $\kappa=8.41$. High-frequency polarizability of tubulin: $\alpha=2.1\times 10^{-33}\,\rm C\,m^{2}/V.$ Sources include both theoretical predictions and experimental results across cryogenic and biological regimes. We also note the finding of electrical oscillations of bundles of microtubules in the brain [37].
@@
Inner lumen diameter 14141515 nm
@@
Typical microtubule length (neuronal) 11100100 µm Varies with cell type[33]
Cryptochrome decoherence time $10^{-5}$ s Predicted, bird magnetoreception [89]
Photosystem I coherence time 300300800800 fs [90]
Light harvesting complex (FMO) coherence >300>300 fs Room temp evidence [90]
@@
Soliton propagation speed (microtubules) 222020 m/s Nonlinear excitations [33]
Time for soliton to travel 1 µm 5050500500 ns $\tau=\frac{1\,\mu m}{v}$
@@
These values are 2233 orders of magnitude smaller than $k_{B}T$, implying that such weak fields alone are insufficient to induce dipole state transitions in thermal equilibrium unless enhanced local fields are used such as those created by defects or nanostructures at a conducting surface. Even without such assistance (which can increase the field strength by several orders of magnitude locally), in the structured environment of the microtubule interior—particularly when interacting with ordered water and neighbouring dipoles—these couplings may contribute to the initiation or modulation of coherent quantum dynamics. In this sense, field-driven transitions may act as subtle biasing agents rather than direct triggers of quantum state evolution.
@@
$\frac{\Delta n}{\Delta c}$ $=(2.0\pm 0.5)\times 10^{-3}\,\text{ml/mg}\,$,
$\frac{\Delta\varepsilon}{\Delta c}$ $=(0.5\pm 0.1)\times 10^{-3}\,\text{ml/mg}\,$, (50)
@@
By varying the plasmonic path length (200–800 nm), it should be possible to measure decoherence times and assess the persistence of entanglement post-interaction, thereby testing the viability of protein-based “bioqubits” in biological quantum computing contexts (see Table 2).